Bowie Sporting Goods manufactures sleeping bags. The manufacturing standards per sleeping bag, based on 5,000 sleeping bags per month, are as follows:

Direct material of 4.50 yards at $5.75 per yard

Direct labor of 2.50 hours at $18.00 per hour

Overhead applied per sleeping bag at $18.00

In the month of April, the company actually produced 4,900 sleeping bags using 24,200 yards of material at a cost of $5.30 per yard. The labor used was 11,500 hours at an average rate of $19.50 per hour. The actual overhead spending was $96,200.

Determine the materials price variance and round to the nearest whole dollar. Enter a favorable variance as a negative number. Enter an unfavorable variance as a positive number.
Answer #1

Materials Price Variance

= (SP-AP) * AQ

= (5.75 - 5.30) * 24200

= 0.45 * 24200

= - 10,890 Favourable(minus)